# Lower bounds for the eigenvalues of the Hodge Laplacian on certain non-convex domains

### v1

@@ Abstract.
We establish geometric lower bounds for the smallest positive eigenvalue of the Hodge Laplacian in the class of non-convex domains given by Euclidean annular regions with a convex outer boundary and a spherical inner boundary. These bounds are then extended to convex domains with multiple holes, where we derive lower bounds for certain higher order exact eigenvalues, and under additional geometric assumptions, also for the smallest positive eigenvalue. For $1$-forms on compact manifolds with boundary, we provide a general lower bound on the smallest exact eigenvalue – corresponding to the first positive Neumann eigenvalue – which, in certain respects, is better than the classical Cheeger inequality. Furthermore, we emphasise the necessity of the contact radius in the lower bounds of the main results. Our proofs employ local-to-global arguments via an explicit isomorphism between Cěch cohomology and de Rham cohomology to obtain Poincaré-type inequalities with explicit geometric dependence, and utilise certain generalised versions of the Cheeger–McGowan gluing lemma.
@@
Hodge Laplacian, eigenvalue bounds, differential forms, annular domains, Cěch cohomology, gluing lemma
@@ 1. Introduction
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{c}}{D}\right)^{2n^{2}-n-2}\min\left\{1,\left(\frac{R_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\}.$
@@
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{c}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\}.$
Here, $D$ and $R_{c}$ denote the diameter and the contact radius of $\mathfrak{D}$, respectively, and $R_{h}$ the radius of the hole.
@@
To prove Theorem 1.1, we adapt the methods of Chanillo and Treves [CT97] – employing tools from Cěch–de Rham theory – in the scenario where the inner hole is sufficiently large compared to the contact radius; while for the case where the hole is much smaller than the contact radius, we use a combination of the bounds from the former case and the spectral convergence of Anne and Colbois [AC93], with the help of a useful gluing lemma by McGowan [McG93]. See Section 3.
@@
Let $n\geq 2$ and $p\in\{1,\dots,n-1\}$. There exists a family $(\mathcal{A}_{\epsilon}^{p})_{{\epsilon}\in(0,\frac{1}{2})}$ of Euclidean annular $C^{1}$-domains with convex outer boundary, such that $\lim_{{\epsilon}\to 0}\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathcal{A}_{\epsilon}^{p})=0$, with the diameter of the domain and the radius of the hole uniformly bounded, and the contact radius tending to zero. It is also implied that the result holds for the smallest positive Neumann eigenvalue for functions, since $\mu_{1}=\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}$.
@@
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{\hat{R}_{c}}{D}\right)^{2n^{2}-n-4+2{\mathfrak{h}}}\min\left\{1,\left(\frac{\hat{R}_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\}.$
@@
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{\hat{R}_{c}R_{h}^{n-1}}{D^{n}}\right)^{{\mathfrak{h}}-1}\left(\frac{\hat{R}_{c}}{D}\right)^{n-2}\min\left\{1,\left(\frac{\hat{R}_{c}}{D}\right)^{3n+2}\left(\frac{\hat{R}_{c}}{R_{h}}\right)^{7n-5}\right\}.$
Although we use the Voronoi partition in the above theorem so as to have $\hat{R}_{c}$ defined independent of the partition, with an optimal contribution to the bounds, the lower bounds also hold if we consider any partition of $\mathfrak{D}$ into $\mathfrak{h}$} many pieces, each of which is a convex domain with a hole in the interior, and modify the definition of $\hat{R}_{c}$ appropriately. Further, similar lower bounds hold for a larger class of domains than those satisfying the above version of the geometric assumption. See Remark 5.4.
@@
We have stated Theorem 1.6 for convex domains with holes of identical size, for simplicity. Analogous lower bounds in the generic setting of convex domains with multiple holes of varying size are presented in Theorem 5.3. For $p\in\{2,\dots,n-1\}$, we further highlight the necessity of the partition parameter $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}$ (see Section 5) in the lower bounds by constructing a family of convex domains with multiple holes (with the diameter of the domain and the radius of the largest hole uniformly bounded), satisfying the hypothesis of Theorem 5.3, having arbitrarily small eigenvalues as $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\to 0$. See Theorem 5.5.
@@ 3. Lower bounds for the smallest positive eigenvalue of convex domains with a hole
In this section, we prove Theorem 1.1 and Theorem 1.3. Let $\mathfrak{D}={\widetilde{\mathfrak{D}}}\setminus\overline{\mathbb{B}_{R_{h}}}$ be an annular domain in $\mathbb{R}^{n}$, where $\widetilde{\mathfrak{D}}$} is a convex domain in $\mathbb{R}^{n}$ and $\mathbb{B}_{R_{h}}\subsetneq{\widetilde{\mathfrak{D}}}$ is a Euclidean $n$-ball of radius $R_{h}$. Let $R_{c}$ and $D$ be the contact radius and the diameter of $\mathfrak{D}$, and $\mathbb{S}_{R_{h}}=\partial\mathbb{B}_{R_{h}}$ denote the round sphere of radius $R_{h}$. Fix $n\geq 2$ and $p\in\{1,\dots,n-1\}$. We divide the proof into two cases– when $R_{h}\geq\frac{R_{c}}{2}$, and when $R_{h}<\frac{R_{c}}{2}$.
@@
where $\tilde{\alpha}$ is an exact form with $i_{\mathpzc{n}}\alpha=0$ and $\tilde{h}$ is a harmonic form satisfying the absolute boundary conditions (hence closed and coclosed). As $\tilde{\omega}$ is closed, we have that $d\delta\tilde{\alpha}=0$, which, using Green’s identity, implies
@@
$\langle\hskip-3.0pt\langle{\omega},{\omega}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$ $=\langle\hskip-3.0pt\langle{\delta\hat{\omega}},{d\tilde{\beta}+\tilde{h}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$
@@
Let us proceed with the proof of Proposition 3.2. We utilise the following estimate for convex domains, by applying it to sufficiently small pieces of the annular domain that are convexified with a control on the volumes and the diameters. We then use these local estimates to deduce a global Poincaré-type inequality, using an explicit isomorphism between Cěch and de Rham cohomology.
@@
Consider the family of constants $(\eta_{I})_{I\in S_{p}}$ as a Čech cochain with respect to the good cover $\{U_{i}\}_{i=1}^{N}$. In view of (3.10), it is a Cěch cocycle. It then follows from (3.7) that the Cěch cohomology class of this cocycle is same as the de Rham cohomology class of $\omega$ (see [BT82, Propostion 9.5]). As $\omega$ is exact, it further means that $(\eta_{I})_{I\in S_{p}}$ is in fact a coboundary. Using this and proceeding as in [CT97, Section 2], we get that
@@
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n(2n-1)}\min\left\{1,\left(\frac{R_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\}.$
@@
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\},$
@@ 4. Construction of small eigenvalues on annular regions
Clearly, the diameter of $\mathcal{A}^{p}_{\epsilon}$ and the radius of the hole are bounded independent of $\epsilon$}. Note that $\{\mathcal{A}^{p}_{\epsilon}\}_{{\epsilon}>0}$ are homotopic to $\mathbb{S}^{n-1}$} and have the cohomology
@@ 5. Eigenvalue lower bounds for convex domains with multiple holes
where $R_{c}$ is the contact radius defined earlier for annular domains. Let $k_{p}$ denote the number of non-empty intersections of order $p$ in the Laguerre-Voronoi partition of $\mathfrak{D}$, and, $\hat{r}_{h}$ and $\hat{R}_{h}$ denote the smallest and the largest of the radii of the holes.
@@
    $\mu^{{(p)}}_{{1+k_{p},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}+n-6}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(n-3)}}}\right\}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
@@
where $r_{c}$ is the distance between the outer boundary and a hole closest to it and $d_{h}$ is the shortest of the distances between any two holes. Further, if $\hat{d}_{h}$ denotes the largest of the distances between any two holes, an elementary computation shows that the partition parameter satisfies
@@
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-4+2{\mathfrak{h}}}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
@@
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1}}{D^{n}}\right)^{{\mathfrak{h}}-1}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\}.$
@@
Note that Hypothesis (H) is always true when $\mathfrak{D}$ has at most two holes. Furthermore, Theorem 5.3 can be adapted to give similar lower bounds for a larger class of domains than those satisfying this hypothesis, for instance, when the cells of the Laguerre-Voronoi partition may be arranged into a finite nested sequence, with the subset of the cells from each inner sequence satisfying Hypothesis (H) and the set of unions of the cells in each inner sequence themselves satisfying an analogue of Hypothesis (H).
@@
Let $n\geq 2$ and $p\in\{1,\dots,n-1\}$. Given $\mathfrak{h}$$\in\mathbb{N}$, there exists a family of Euclidean $C^{1}$-domains with convex outer boundary and $\mathfrak{h}$} many holes, satisfying Hypothesis (H) and having arbitrarily small exact eigenvalue, with the diameter of the domain and the radius of the largest hole uniformly bounded, and the partition parameter tending to zero.
@@
Then we have from Theorem 1.1 that, for $n\geq 3$ and $p\in\{2,\dots n-1\}$,
@@
$\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}+n-6}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(n-3)}}}\right\}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
@@
So, it will suffice to construct such a pair with the above quotient of the (squares of the) $L^{2}$-norms bounded below by appropriate terms. Given a $\phi$, we can locally construct $\psi$ using the eigenvalues of the pieces $\{U_{i}\}$ and their intersections. We then use a partition of unity for the local-to-global transition employing the Cěch–de Rham formalism, which enforces a choice of $\phi$ to be made in the proof, as we shall see.
@@
Figure 8. Cěch–de Rham formalism
@@
with $\check{C}^{m}(\{U_{i}\},\mathbb{R})$ denoting the space of locally constant functions on non-trivial $m$-intersections of $\{U_{i}\}$. This sequence is a differential complex, and its homology is called the Cěch cohomology of the cover $\{U_{i}\}_{i=1}^{k_{0}}$. See [BT82, Chapter 8] for a detailed description of the Čech-de Rham formalism.

### v2

@@ Abstract.
We establish geometric lower bounds for the smallest positive eigenvalue of the Hodge Laplacian in the class of non-convex domains given by Euclidean annular regions with a convex outer boundary and a spherical inner boundary. These bounds are then extended to convex domains with multiple holes, where we derive lower bounds for certain higher order exact eigenvalues, and under additional geometric assumptions, also for the smallest positive eigenvalue. For $1$-forms on compact manifolds with boundary, we provide a general lower bound on the smallest exact eigenvalue – corresponding to the first positive Neumann eigenvalue – which, in certain respects, is better than the classical Cheeger inequality. Furthermore, we emphasise the necessity of the contact radius in the lower bounds of the main results. Our proofs employ local-to-global arguments via an explicit isomorphism between Čech cohomology and de Rham cohomology to obtain Poincaré-type inequalities with explicit geometric dependence, and utilise certain generalised versions of the Cheeger–McGowan gluing lemma.
@@
Hodge Laplacian, eigenvalue bounds, differential forms, annular domains, Čech cohomology, gluing lemma
@@ 1. Introduction
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{c}}{D}\right)^{5n^{2}-n-5}.$
@@
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{c}}{D}\right)^{11n-5}.$
Here, $D$ and $R_{c}$ denote the diameter and the contact radius of $\mathfrak{D}$, respectively.
@@
In particular, we have new lower bounds for the Neumann eigenvalues of annular domains. Note that, for planar domains, the spectrum for $1$-forms consists solely of the Neumann and the Dirichlet eigenvalues for functions, and understanding these would already give a complete picture of the full spectrum of the Hodge Laplacian for $p\in\{0,1,2\}$.
To prove Theorem 1.1, we adapt the methods of Chanillo and Treves [CT97] – employing tools from Čech–de Rham theory – in the scenario where the inner hole is sufficiently large compared to the contact radius; while for the case where the hole is much smaller than the contact radius, we use a combination of the bounds from the former case and the spectral convergence of Anné and Colbois [AC93], with the help of a useful gluing lemma by McGowan [McG93]. See Section 3.
@@
Let $n\geq 2$ and $p\in\{1,\dots,n-1\}$. There exists a family $(\mathcal{A}_{\epsilon}^{p})_{{\epsilon}\in(0,\frac{1}{2})}$ of Euclidean annular $C^{1}$-domains with convex outer boundary, such that $\lim_{{\epsilon}\to 0}\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathcal{A}_{\epsilon}^{p})=0$, with the diameter uniformly bounded and the contact radius tending to zero. It is also implied that the result holds for the smallest positive Neumann eigenvalue for functions, since $\mu_{1}=\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}$.
@@
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{\hat{R}_{c}}{D}\right)^{5n^{2}-n-7+2{\mathfrak{h}}}.$
@@
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{\hat{R}_{c}R_{h}^{n-1}}{D^{n}}\right)^{{\mathfrak{h}}-1}\left(\frac{\hat{R}_{c}}{D}\right)^{11n-5}.$
A simple example of a domain satisfying the above hypothesis is as follows.
Figure 3. Example I
Although we use the Voronoi partition in the above theorem so as to have $\hat{R}_{c}$ defined independent of the partition, with an optimal contribution to the bounds, the lower bounds also hold if we consider any partition of $\mathfrak{D}$ into $\mathfrak{h}$} many pieces, each of which is a convex domain with a hole in the interior, and modify the definition of $\hat{R}_{c}$ appropriately. This also allows, for example, domains of the following type whose Voronoi partitions themselves may not satisfy our hypothesis unlike Example I.
Figure 4. Example II
Further, similar lower bounds can be obtained for a larger class of domains than those satisfying the above versions of the geometric assumption. See Remark 5.4. This allows in addition, the following domain, for instance, which is a union of the above two examples (see also Figure 11).
Figure 5. Example III
The following, however, is a non-example. For this domain, no partition with convex outer boundaries satisfies our hypothesis, nor can we apply the extension provided by Remark 5.4.
Figure 6. Non-example
@@
We have stated Theorem 1.6 for convex domains with holes of identical size, for simplicity. Analogous lower bounds in the generic setting of convex domains with multiple holes of varying size are presented in Theorem 5.3. For $p\in\{2,\dots,n-1\}$, we further highlight the necessity of the partition parameter $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}$ (see Section 5) in the lower bounds by constructing a family of convex domains with multiple holes (with the diameter of the domain), satisfying the hypothesis of Theorem 5.3, having arbitrarily small eigenvalues as $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\to 0$. See Theorem 5.5.
@@ 3. Lower bounds for the smallest positive eigenvalue of convex domains with a hole
In this section, we prove Theorem 1.1 and Theorem 1.3. Before that, we give an overview of the proof of Theorem 1.1. We consider the two regimes—when the radius of the hole $R_{h}$ is large, or small, compared to the contact radius $R_{c}$. In the former case, we begin with an eigenform $\omega$ of the first exact eigenvalue $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$ and would like to find a primitive of it which satisfies a Poincaré-type inequality with a geometrically explicit constant, (the square of) whose inverse would serve as a lower bound for $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$. Such an inequality is known for convex domains from the work of Chanillo-Treves [CT97, Lemma 2.1], and the geometric dependence of the constant there is suitable to bring out the role of $R_{c}$ by adapting it to our setting. We consider an open cover of our annular domain by dividing it into sufficiently small pieces that are preimages of tiny patches on the inner spherical boundary, under the orthogonal projection map of $\mathbb{R}^{n}$ onto it. The boundary of each piece has a convex part, coming from the outer boundary of the initial domain, and a non-convex part that is a remanent of the inner boundary, in addition to the flat part of the boundary connecting these two. Now, we convexify each piece near the non-convex part using a quasi-isometry (see Figure 8), with explicit control on the distortion required to do so. We then apply [CT97, Lemma 2.1] to these convex domains that approximate our pieces to obtain local Poincaré-type inequalities for the pieces themselves, involving the restrictions of the eigenform $\omega$. However, the primitives thus obtained on these pieces do not necessarily agree on the intersections for them to be glued together. In order to obtain a global primitive, we use an appropriate partition of unity to first construct an approximate global $p$-form $\nu$ from the primitives on the pieces. We then consider the difference $\omega-d\nu$, and recursively update $\nu$ in $p$ steps to eventually obtain a more manageable expression for $\omega-d\nu$ which involves the first derivatives of the partition of unity and certain constant functions on larger versions of the pieces (see Lemma 3.5 for a precise description). These constants, when viewed as a Čech cochain for a suitable open cover of the domain, are in fact a Čech cocycle. Furthermore, the exactness of $\omega$ implies that this system of constants is in fact a Čech coboundary. Using an explicit isomorphism between the Čech cocyles and the closed forms representing the de Rham cohomology [BT82, Proposition 9.8], this allows us to further simplify the difference $\omega-d\nu$ and finally construct a primitive $\eta$ of $\omega$. Throughout this process, we keep track of the estimates and obtain a global Poincaré-type inequality as intended. In the latter regime of small holes, we proceed as follows. A spectral convergence result of Anné–Colbois [AC93, Theorem 1.1] for perforated domains with shrinking spherical holes gives explicit lower bounds when the domains are annuli whose inner and outer radii are chosen appropriately. When we have a generic domain with convex outer boundary and a spherical hole, we consider a $2$-cover comprising an inner annulus of suitable dimensions and an outer annular region whose hole now has sufficiently large radius for the lower bounds from the earlier regime to be applicable. Finally, we apply to this $2$-cover a gluing lemma of Cheeger–McGowan [McG93, Lemma 2.3] that, under certain topological assumptions which hold in our setting, gives lower bounds for the exact $p$-spectrum of a domain in terms of the lower bounds on the exact $p$-spectrum of the two pieces and the exact (p1)(p-1)-spectrum of their intersection. This gluing procedure gives lower bounds on $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$ for $p\in\{2,\dots,n-1\}$. For the case $p=1$ (which also corresponds to the Neumann eigenvalue for functions), the role of this gluing lemma is replaced by our Theorem 1.3. We now proceed to the proof of Theorem 1.1.
Let $\mathfrak{D}={\widetilde{\mathfrak{D}}}\setminus\overline{\mathbb{B}_{R_{h}}}$ be an annular domain in $\mathbb{R}^{n}$, where $\widetilde{\mathfrak{D}}$} is a convex domain in $\mathbb{R}^{n}$ and $\mathbb{B}_{R_{h}}\subsetneq{\widetilde{\mathfrak{D}}}$ is a Euclidean $n$-ball of radius $R_{h}$. Let $R_{c}$ and $D$ be the contact radius and the diameter of $\mathfrak{D}$, and $\mathbb{S}_{R_{h}}=\partial\mathbb{B}_{R_{h}}$ denote the round sphere of radius $R_{h}$. Fix $n\geq 2$ and $p\in\{1,\dots,n-1\}$. We divide the proof into two cases– when $R_{h}\geq\frac{R_{c}}{2}$, and when $R_{h}<\frac{R_{c}}{2}$.
@@
where $\tilde{\alpha}$ is an exact form with $i_{\mathpzc{n}}\tilde{\alpha}=0$ and $\tilde{h}$ is a harmonic form satisfying the absolute boundary conditions (hence closed and coclosed). As $\tilde{\omega}$ is closed, we have that $d\delta\tilde{\alpha}=0$, which, using Green’s identity, implies
@@
$\langle\hskip-3.0pt\langle{\tilde{\omega}},{\tilde{\omega}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$ $=\langle\hskip-3.0pt\langle{\delta\hat{\omega}},{d\tilde{\beta}+\tilde{h}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$
@@
Let us proceed with the proof of Proposition 3.2. We utilise the following estimate for convex domains, by applying it to sufficiently small pieces of the annular domain that are convexified with a control on the volumes and the diameters. We then use these local estimates to deduce a global Poincaré-type inequality, using an explicit isomorphism between Čech and de Rham cohomology.
@@
Consider the family of constants $(\eta_{I})_{I\in S_{p}}$ as a Čech cochain with respect to the good cover $\{U_{i}\}_{i=1}^{N}$. In view of (3.10), it is a Čech cocycle. It then follows from (3.7) that the Čech cohomology class of this cocycle is same as the de Rham cohomology class of $\omega$ (see [BT82, Propostion 9.5]). As $\omega$ is exact, it further means that $(\eta_{I})_{I\in S_{p}}$ is in fact a coboundary. Using this and proceeding as in [CT97, Section 2], we get that
@@
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n(2n-1)}\min\left\{1,\left(\frac{R_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\},$ (3.16)
giving the required lower bound, using that $\hat{R}_{h}\leq D$.
@@
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\},$ (3.17)
@@ 4. Construction of small eigenvalues on annular regions
Clearly, the diameter of $\mathcal{A}^{p}_{\epsilon}$ is bounded independent of $\epsilon$}. Note that $\{\mathcal{A}^{p}_{\epsilon}\}_{{\epsilon}>0}$ are homotopic to $\mathbb{S}^{n-1}$} and have the cohomology
@@ 5. Eigenvalue lower bounds for convex domains with multiple holes
where $R_{c}$ is the contact radius defined earlier for annular domains. Let $k_{p}$ denote the number of non-empty intersections of order $p$ in the Laguerre-Voronoi partition of $\mathfrak{D}$.
@@
    $\mu^{{(p)}}_{{1+k_{p},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{5n^{2}+n-9}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(n-3)}}}\right\}.$
@@
    $\mu^{{(1)}}_{{1+k_{1},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{11n-7}.$
@@
where $r_{c}$ is the distance between the outer boundary and a hole closest to it and $d_{h}$ is the shortest of the distances between any two holes. Further, if $\hat{d}_{h}$ denotes the largest of the distances between any two holes, and $\hat{r}_{h}$, $\hat{R}_{h}$ denote the smallest and the largest of the radii of the holes, an elementary computation shows that the partition parameter satisfies
@@
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{5n^{2}-n-7+2{\mathfrak{h}}}.$
@@
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1}}{D^{n}}\right)^{{\mathfrak{h}}-1}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{11n-7}.$
@@
Note that Hypothesis (H) is always true when $\mathfrak{D}$ has at most two holes. Furthermore, Theorem 5.3 can be adapted to give similar lower bounds for a larger class of domains than those satisfying this hypothesis, for instance, when the cells of the Laguerre-Voronoi partition may be arranged into a finite nested sequence, with the subset of the cells from each inner sequence satisfying Hypothesis (H) and the set of unions of the cells in each inner sequence themselves satisfying an analogue of Hypothesis (H). This allows, for example, the domain depicted in Figure 5. Another simple example is the following domain with four holes, which does not directly satisfy Hypothesis (H) but we can apply this remark to get lower bounds on $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$.
Figure 11. Example IV
@@
Let $n\geq 2$ and $p\in\{1,\dots,n-1\}$. Given $\mathfrak{h}$$\in\mathbb{N}$, there exists a family of Euclidean $C^{1}$-domains with convex outer boundary and $\mathfrak{h}$} many holes, satisfying Hypothesis (H) and having arbitrarily small exact eigenvalue, with the diameter of the domain uniformly bounded and the partition parameter tending to zero.
@@
To prove Theorem 5.1, we consider an open cover of $\mathfrak{D}$ given by the Laguerre-Voronoi partition and apply certain generalised versions of McGowan’s gluing lemma (see Lemma 5.6 and Appendix) in combination with the lower bounds (3.16), (3.17) of Theorem 1.1 applied to the cells.
@@
Then we have from (3.16) that, for $n\geq 3$ and $p\in\{2,\dots n-1\}$,
@@
$\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}+n-6}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(n-3)}}}\right\}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\},$
giving the required lower bound, using that $\hat{R}_{h}\leq D$. ∎
In order to obtain lower bounds for the first exact eigenvalue, we instead use directly Lemma 3.7 by starting with a cell in the open cover and adding cells inductively such that the convexity of the outer boundary of the intermediate domain is preserved at each step (there exists such an ordering guaranteed by Hypothesis (H)), and invoking once again the lower bounds (3.16), (3.17) of Theorem 1.1 for the cells being added. This procedure introduces the dependence of the exponents of the ratio $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}/D$ on the number of holes. As before, for $p=1$, the gluing lemma needs to be replaced by Theorem 1.3 and we adapt our approach accordingly, as described in the proof below.
@@
So, it will suffice to construct such a pair with the above quotient of the (squares of the) $L^{2}$-norms bounded below by appropriate terms. Given a $\phi$, we can locally construct $\psi$ using the eigenvalues of the pieces $\{U_{i}\}$ and their intersections. We then use a partition of unity for the local-to-global transition employing the Čech–de Rham formalism, which enforces a choice of $\phi$ to be made in the proof, as we shall see.
@@
Figure 13. Čech–de Rham formalism
@@
with $\check{C}^{m}(\{U_{i}\},\mathbb{R})$ denoting the space of locally constant functions on non-trivial $m$-intersections of $\{U_{i}\}$. This sequence is a differential complex, and its homology is called the Čech cohomology of the cover $\{U_{i}\}_{i=1}^{k_{0}}$. See [BT82, Chapter 8] for a detailed description of the Čech-de Rham formalism.